# Robust Bernoulli mixture models for credit portfolio risk

### v1

@@ 5.1 Simulation study
    using (5.2) and (3.8). Since the Clayton copula family is increasing in $\theta$ with respect to the pointwise order and since Clayton copulas are SI for a non-negative parameter, we obtain the following sharp lower and upper default integral bounds:
    $\underline{G}^{Cl}_{n}(s)$ $:=\bigvee_{\theta_{n}\in I}G_{D_{n}^{\theta_{n}},Z}(s)=G_{D_{n}^{\overline{\theta}},Z}(s)=\pi_{n}-\left(\pi_{n}^{-\overline{\theta}}+(1-s)^{-\overline{\theta}}-1\right)^{-1/\overline{\theta}}$ (5.8)
    $\overline{G}^{Cl}_{n}(s)$ $:=\bigwedge_{\theta_{n}\in I}G_{D_{n}^{\theta_{n}},Z}(s)=G_{D_{n}^{\underline{\theta}},Z}(s)=\pi_{n}-\left(\pi_{n}^{-\underline{\theta}}+(1-s)^{-\underline{\theta}}-1\right)^{-1/\underline{\theta}}\,.$ (5.9)
    The upper right plot in Figure 5.1 illustrates the range of default integral functions in the Clayton threshold model under parameter uncertainty, where, for all $n\in\{1,\ldots,N\}\,,$ the upper (lower) dashed graph corresponds to the lower (upper) default integral bound $\underline{G}^{Cl}_{n}$ ($\overline{G}^{Cl}_{n}$). The solid line indicates the default integral function for $\theta_{n}\approx 0.723$ for which the Clayton copula has the same Kendall’s tau value as the Gaussian copula in (a) obtained from the IRB formula.
@@
    Survival Clayton copula threshold model: We consider the setting in (b), where we now replace the Clayton copula family by the survival Clayton copula family. Noting that Kendall’s tau value and the SI property are invariant under taking the survival copula transformation (3.8), we obtain the lower and upper default integral bounds
    $\underline{G}^{sCl}_{n}(s)$ $:=\bigvee_{\theta_{n}\in I}G_{D_{n}^{\theta_{n}},Z}(s)=G_{D_{n}^{\overline{\theta}},Z}(s)=s-\left((1-\pi_{n})^{-\overline{\theta}}+s^{-\overline{\theta}}-1\right)^{-1/\overline{\theta}}\,,$ (5.10)
    $\overline{G}^{sCl}_{n}(s)$ $:=\bigwedge_{\theta_{n}\in I}G_{D_{n}^{\theta_{n}},Z}(s)=G_{D_{n}^{\underline{\theta}},Z}(s)=s-\left((1-\pi_{n})^{-\underline{\theta}}+s^{-\underline{\theta}}-1\right)^{-1/\underline{\theta}}\,.$ (5.11)
@@
Tables 5.1 and 5.2 highlight the significant impact of accounting for lower tail dependencies in a credit portfolio comprising $N=1.000$ loans, each with a homogeneous default probability of $\pi_{n}=0.02$ for all $n\in\{1,\ldots,N\}\,.$ As underlying convex risk measure, we choose the $\operatorname{AVaR}_{\alpha}$ with levels $\alpha=95\%$ and $\alpha=99\%\,.$ In the first scenario, we consider a deterministic loss given default of $\delta_{n}\equiv 0.1$ (see Table 5.1); in the second scenario, we choose a beta-distributed loss given default $\delta_{n}$ with expectation $\mathbb{E}[\delta_{n}]=0.1$ and volatility $\sqrt{\mathrm{Var}(\delta_{n})}=0.15$ (see Table 5.2). In the settings of the default models analysed in Example 5.1, our simulations demonstrate that accurately capturing tail dependencies is more critical than accounting for parameter uncertainty within the individual models. This is evidenced by the non-overlapping uncertainty intervals observed in the Gaussian, Clayton, and survival Clayton copula settings. The lower bound of the uncertainty interval in the survival Clayton copula setting is only marginally higher than the risk $\operatorname{AVaR}_{\alpha}^{\perp}:=\operatorname{AVaR}_{\alpha}(L^{\perp})$ of the credit portfolio with independent default events. This can be attributed to the fact that the lower tail dependencies in the survival Clayton copula are nearly independent (as evident from the lower left corner of the sample generated from the survival Clayton copula in Figure 3.1). In contrast, the Clayton copula setting captures lower tails with a notably strong positive dependence, effectively accounting for the pronounced co-movements in extreme events (see the lower left corner of the sample from the Clayton copula in Figure 3.1). As a consequence, the upper bound for the Clayton copula is about two third of the worst-case comonotonic bound $\operatorname{AVaR}_{\alpha}^{c}:=\operatorname{AVaR}_{\alpha}(L^{c})\,.$ The tail dependencies in the Gaussian setting lie somewhere in between, suggesting that portfolio risks modeled using the Gaussian model might initially appear to provide a realistic approach. However, as previously noted, Gaussian copulas typically underestimate tail dependencies in financial applications. Consequently, the Clayton copula setting may offer a more accurate reflection of the associated risks.
@@
To find a balance between the Gaussian and the Clayton copula settings, we introduce a class of distributions whose tail dependencies lie between those of the specified Gaussian and Clayton copula. More precisely, we consider a class siBMMs, where the default integral functions are generated by a combination of Gaussian and Clayton copula threshold models discussed before.
Figure 5.1: Default integral functions $G_{D_{n},Z}$ of the threshold model classes considered in Examples 5.1 and 5.2 under parameter and model uncertainty: Gaussian copula model (green), Clayton copula model (blue), and survival Clayton copula model (red). The solid lines are obtained from asset correlations estimated by the IRB formula, where the dashed lines correspond to the asset correlation bounds of the IRB formula (5.3); see (5.4)–(5.11) for a specification of these functions.
Example 5.2 (Bernoulli mixture models under model uncertainty).
@@
We consider a class of siBMMs specified by the default probabilities $\pi_{n}=0.02$ for all $n$ and by the conditional default probability functions $p_{D_{n}}$ associated with the class $\mathfrak{c}(\{G_{n}^{Ga},G_{n}^{Cl}\})$ (defined in (4.9)) of default integral functions. Here $G_{n}^{Ga}$ is the default integral function in (5.4) associated with the Gaussian copula with parameter $\rho_{n}\approx 0.165$ implied by the IRB formula. The function $G_{n}^{Cl}$ is the default integral function in (5.7) associated with the Clayton copula with parameter $\theta_{n}\approx 0.723\,,$ which accounts for tail dependencies. The lower right-hand plot in Figure 5.1 illustrates the class $\mathfrak{c}(\{G_{n}^{Ga},G_{n}^{Cl}\})$ of default integral functions. We observe that the derivative of the function $G_{n}^{Cl}$ (blue graph) is larger than that of $G_{n}^{Ga}$ (green plot) for $s$ close to $1\,.$ This implies that the probability of default increases for large values of the common risk factor $Z$, leading to a higher likelihood of simultaneous defaults
@@
Table 5.1: Lower and upper bounds for the AVaR of the percentage losses in the robust credit risk models under parameter uncertainty and model uncertainty as specified in Examples 5.1 and 5.2. The number of loans is chosen as $N=1.000\,,$ the loss given default is $\delta_{n}\equiv 0.1$ and the exposure at default as fraction of the total amount is $e_{n}=1/N$ for all $n\in\{1,\ldots,N\}\,.$ The numbers are obtained by Monte Carlo simulations with a sample size of $1.000.000\,.$
We apply Corollary 4.7, to determine the risk bounds for the AVaR in the hybrid ’Gauss-Clayton’ setting specified in Example 5.2. The results are illustrated in Tables 5.1 and 5.2 for the two different scenarios of loss given default as discussed previously. The risk bounds capture model uncertainty for the class of default integral functions lying between the Gaussian and Clayton case. Figure 5.1 illustrates the classes of default integral functions considered in the four cases in Examples 5.1 and 5.2.
@@
Table 5.2: Lower and upper bound for the AVaR of the percentage losses in the robust credit risk models under parameter uncertainty and model uncertainty as specified in Examples 5.1 and 5.2. The number of loans is chosen as $N=1.000\,,$ the loss given default $\delta_{n}$ is $B(0.3,2.7)$-distributed, and the exposure at default as fraction of the total amount is $e_{n}=1/N$ for all $n\in\{1,\ldots,N\}\,.$ The numbers are obtained by Monte Carlo simulations with a sample size of $1.000.000\,.$
@@ 5.2 Real data example
Table 5.3: Borrowing countries, credit rating 666according to S&P, Fitch or Moody’s ratings, https://tradingeconomics.com/, or obtained by regressing OECD ratings on S&P ratings if ratings were not available and amount outstanding (in million USD) as of 2022 (Source: IDB (2022)). $\pi_{n}$ denotes the default probability and $e_{n}$ is the relative amount of the exposure in the total portfolio. The last two pairs of columns refer the asset correlation uncertainty intervals $I_{n}=[\underline{\rho}_{n},\overline{\rho}_{n}]$ in the Gaussian model (Example 5.1(a)) and the parameter uncertainty interval $I_{n}=[\underline{\theta}_{n},\overline{\theta}_{n}]$ in the Clayton models (Examples 5.1(b) and (c)), where we set $\underline{\rho}_{n}:=\rho_{n}-5\%$ and $\overline{\rho}_{n}:=\rho_{n}+5\%$ with the asset correlation $\rho_{n}$ obtained from the IRB formula (5.3) depending on $\pi_{n}$ and the bounds chosen as $0.11$ and $0.27\,.$
66footnotetext: According to S&P, Fitch or Moody’s ratings, https://tradingeconomics.com/, or obtained by regressing OECD ratings on S&P ratings if ratings were not available.
@@
Table 5.4: Lower and upper bounds for the AVaR of the percentage losses for the real-data example discussed in Section 5.2. The underlying robust credit risk models incorporate parameter uncertainty and model uncertainty in a Gaussian, Clayton and survival Clayton copula setting as in Examples 5.1 and 5.2, see the caption of Table 6 for the respective parameters. The loss given default in this scenario is $\delta_{n}\equiv 0.1$ and the numbers are obtained by Monte Carlo simulations with a sample size of $1.000.000\,.$
@@
$\alpha=99$ 11.19% 11.22% 11.27% 11.36% 11.16% 11.17% 11.21% 11.32% 11.18% 11.55%
Table 5.5: Lower and upper bounds for the AVaR of the percentage losses for the real-data example discussed in Section 5.2. The underlying robust credit risk models incorporate parameter uncertainty and model uncertainty in a Gaussian, Clayton and survival Clayton copula setting as in Examples 5.1 and 5.2, see the caption of Table 6 for the respective parameters. The loss given default in this scenario is $\delta_{n}\sim B(0.3,2.7)$-distributed and the numbers are obtained by Monte Carlo simulations with a sample size of $1.000.000\,.$
@@
While in the case of a homogeneous and rather large credit risk portfolio (Tables 5.1 and 5.2), the dependencies between the default events are crucial, we find that in the rather small (in terms of number of borrowers) and very heterogeneous credit portfolio of IDB, the AVaR bounds are much more comparable (Tables 5.4 and 5.5). The reason is that for such a small and concentrated (w.r.t. single borrowers) portfolio, systematic risk reflected by common dependence on the risk factor is less relevant whereas a diversified idiosyncratic risk plays a much more important role, compare Lütkebohmert et al. (2023) and Lütkebohmert and Sester (2024). Moreover, we observe that the distribution of the losses given default has a strong influence on the total portfolio risk which is due to the fact that tail events result from a combination of bad realizations of the risk factor and high values for the loss given default variable. Thus, AVaR increases with the variance of loss given default.
@@ Conclusion
In this paper, we established a convex ordering result for credit portfolio losses in stochastically increasing Bernoulli mixture models based on transparent and interpretable assumptions on the conditional default probabilities. For threshold models, we provided a related ordering result based on a pointwise comparison of bivariate copulas. As a consequence, we obtained several robustness results for credit portfolio losses. Without assuming monotonicity of the conditional default probabilities in the common factor, we derived an upper bound on portfolio losses that improves the standard comonotonic bound. A detailed simulation study and a real-data example illustrate the relevance and effectiveness of our approach. Our methods are especially applicable to standard industry models of credit portfolio risk, such as CreditMetrics or KMV Portfolio Manager, and allow for the incorporation of tail dependencies and model uncertainty into these models. In this way, our paper provides tractable and practically relevant methods for the assessment of credit portfolio risk under mild structural assumptions.
@@ Acknowledgements
The first author gratefully acknowledges the support of the Austrian Science Fund (FWF) project P 36155-N ReDim: Quantifying Dependence via Dimension Reduction and the support of the WISS 2025 project ’IDA-lab Salzburg’ (20204-WISS/225/197-2019 and 20102-F1901166-KZP).

### v2

@@ 4.1 Simulation study
    using (4.1). Since the Clayton copula family is increasing in $\theta$ with respect to the pointwise order and since Clayton copulas are SI for a non-negative parameter, we obtain the following sharp lower and upper default integral bounds:
    $\overline{G}^{Cl}_{n}(s)$ $:=\underline{\mathop{co}}\big((G_{D_{n}^{\theta_{n}},Z})_{\rho_{n}\in I})\big)(s)=G_{D_{n}^{\overline{\theta}},Z}(s)=\pi_{n}-\left(\pi_{n}^{-\overline{\theta}}+(1-s)^{-\overline{\theta}}-1\right)^{-1/\overline{\theta}},$
    $\underline{G}^{Cl}_{n}(s)$ $:=\overline{\mathop{co}}\left((G_{D_{n}^{\rho_{n}},Z})_{\rho_{n}\in I})\right)(s)=G_{D_{n}^{\underline{\theta}},Z}(s)=\pi_{n}-\big(\pi_{n}^{-\underline{\theta}}+(1-s)^{-\underline{\theta}}-1\big)^{-1/\underline{\theta}}.$
    The upper right plot in Figure 4.1 illustrates the range of default integral functions in the Clayton threshold model under parameter uncertainty, where, for all $n\in\{1,\ldots,N\},$ the lower (upper) dashed graph corresponds to the default integral bound $\overline{G}^{Cl}_{n}$ ($\underline{G}^{Cl}_{n}$). The solid line indicates the default integral function for $\theta_{n}\approx 0.723$ for which the Clayton copula has the same Kendall’s tau value as the Gaussian copula in (a) obtained from the IRB formula.
@@
    Survival Clayton copula threshold model: We consider the setting in (b), where we now replace the Clayton copula family in (4.4) by the survival Clayton copula family. Noting that Kendall’s tau value and the SI property are invariant under taking the survival copula transformation (A.7), we obtain the lower and upper default integral bounds
    $\overline{G}^{sCl}_{n}(s)$ $:=\underline{\mathop{co}}\big((G_{D_{n}^{\theta_{n}},Z})_{\rho_{n}\in I})\big)(s)=G_{D_{n}^{\overline{\theta}},Z}(s)=s-\left((1-\pi_{n})^{-\overline{\theta}}+s^{-\overline{\theta}}-1\right)^{-1/\overline{\theta}},$
    $\underline{G}^{sCl}_{n}(s)$ $:=\overline{\mathop{co}}\big((G_{D_{n}^{\rho_{n}},Z})_{\rho_{n}\in I})\big)(s)=G_{D_{n}^{\underline{\theta}},Z}(s)=s-\left((1-\pi_{n})^{-\underline{\theta}}+s^{-\underline{\theta}}-1\right)^{-1/\underline{\theta}}.$
@@
Tables 4.1 and 4.2 highlight the significant impact of accounting for lower tail dependencies in a credit portfolio comprising $N=1.000$ loans, each with a homogeneous default probability of $\pi_{n}=0.02$ for all $n\in\{1,\ldots,N\}.$ As underlying convex risk measure, we choose the $\operatorname{AVaR}_{\alpha}$ with levels $\alpha=95\%$ and $\alpha=99\%.$ In the first scenario, we consider a deterministic loss given default of $\delta_{n}\equiv 0.1$ (see Table 4.1); in the second scenario, we choose a beta-distributed loss given default $\delta_{n}$ with expectation $\mathbb{E}[\delta_{n}]=0.1$ and volatility $\sqrt{\mathrm{Var}(\delta_{n})}=0.15$ (see Table 4.2). In the settings of the default models analysed in Example 4.1, our simulations demonstrate that accurately capturing tail dependencies is more critical than accounting for parameter uncertainty within the individual models. This is evidenced by the non-overlapping uncertainty intervals observed in the Gaussian, Clayton, and survival Clayton copula settings. The lower bound of the uncertainty interval in the survival Clayton copula setting is only marginally higher than the risk $\operatorname{AVaR}_{\alpha}^{\perp}:=\operatorname{AVaR}_{\alpha}(L^{\perp})$ of the credit portfolio with independent default events. This can be attributed to the fact that the lower tail dependencies in the survival Clayton copula are nearly independent (as evident from the lower left corner of the sample generated from the survival Clayton copula in Figure A.1). In contrast, the Clayton copula setting captures lower tails with a notably strong positive dependence, effectively accounting for the pronounced co-movements in extreme events (see the lower left corner of the sample from the Clayton copula in Figure A.1). As a consequence, the upper bound for the Clayton copula is about two third of the worst-case comonotonic bound $\operatorname{AVaR}_{\alpha}^{c}:=\operatorname{AVaR}_{\alpha}(L^{c}).$ The tail dependencies in the Gaussian setting lie somewhere in between, suggesting that portfolio risks modeled using the Gaussian model might initially appear to provide a realistic approach. However, as previously noted, Gaussian copulas typically underestimate tail dependencies in financial applications. Consequently, the Clayton copula setting may offer a more accurate reflection of the associated risks.
@@
To find a balance between the Gaussian and the Clayton copula settings, we introduce a class of distributions whose tail dependencies lie between those of the specified Gaussian and Clayton copula. More precisely, we consider a class siBMMs, where the default integral functions are generated by a combination of Gaussian and Clayton copula threshold models.
Example 4.2 (siBMMs under model uncertainty).
@@
We consider a class of siBMMs specified by the default probabilities $\pi_{n}=0.02$ for all $n$ and by the conditional default probability functions $p_{D_{n}}$ associated with the class $\mathfrak{c}(\{G_{n}^{Ga},G_{n}^{Cl}\})$ (defined in (3.5)) of default integral functions. Here $G_{n}^{Ga}$ is the default integral function in (4.3) associated with the Gaussian copula with parameter $\rho_{n}\approx 0.165$ implied by the IRB formula. The function $G_{n}^{Cl}$ is the default integral function in (4.4) associated with the Clayton copula with parameter $\theta_{n}\approx 0.723\,$, which accounts for tail dependencies. The lower right-hand plot in Figure 4.1 illustrates the class $\mathfrak{c}(\{G_{n}^{Ga},G_{n}^{Cl}\})$ of default integral functions. We observe that the derivative of the function $G_{n}^{Cl}$ (blue graph) is larger than that of $G_{n}^{Ga}$ (green plot) for $s$ close to $1.$ This implies that the probability of default increases for large values of the common risk factor $Z$, leading to a higher likelihood of simultaneous defaults.
@@
Table 4.1: Lower and upper bounds for the AVaR of the percentage losses in the robust credit risk models under parameter uncertainty and model uncertainty as specified in Examples 4.1 and 4.2. The number of loans is chosen as $N=1.000,$ the loss given default is $\delta_{n}\equiv 0.1$ and the exposure at default as fraction of the total amount is $e_{n}=1/N$ for all $n\in\{1,\ldots,N\}.$ The numbers are obtained by Monte Carlo simulations with a sample size of $1.000.000.$
The risk bounds for the AVaR in the hybrid ’Gauss-Clayton’ setting specified in Example 4.2 are also illustrated in Tables 4.1 and 4.2 for the two different scenarios of loss given default as discussed previously. The risk bounds capture model uncertainty for the class of default integral functions lying between the Gaussian and Clayton case. Recall that Figure 4.1 illustrates the classes of default integral functions considered in Examples 4.1 and 4.2.
@@
Table 4.2: Lower and upper bound for the AVaR of the percentage losses in the robust credit risk models under parameter uncertainty and model uncertainty as specified in Examples 4.1 and 4.2. The number of loans is chosen as $N=1.000,$ the loss given default $\delta_{n}$ is $beta(0.3,2.7)$-distributed, and the exposure at default as fraction of the total amount is $e_{n}=1/N$ for all $n\in\{1,\ldots,N\}.$ The numbers are obtained by Monte Carlo simulations with a sample size of $1.000.000.$
@@ 4.2 Real data example
Table 4.3: Borrowing countries, credit rating 444according to S&P, Fitch or Moody’s ratings, https://tradingeconomics.com/, or obtained by regressing OECD ratings on S&P ratings if ratings were not available and amount outstanding (in million USD) as of 2022 (Source: (IDB, 2022, p.83)). $\pi_{n}$ denotes the default probability and $e_{n}$ is the relative amount of the exposure in the total portfolio. The last two pairs of columns refer the asset correlation uncertainty intervals $I_{n}=[\underline{\rho}_{n},\overline{\rho}_{n}]$ in the Gaussian model (Example 4.1(a)) and the parameter uncertainty interval $I_{n}=[\underline{\theta}_{n},\overline{\theta}_{n}]$ in the Clayton models (Examples 4.1(b) and (c)), where we set $\underline{\rho}_{n}:=\rho_{n}-5\%$ and $\overline{\rho}_{n}:=\rho_{n}+5\%$ with the asset correlation $\rho_{n}$ obtained from the IRB formula (4.2) depending on $\pi_{n}$ and the bounds chosen as $0.11$ and $0.27.$
44footnotetext: According to S&P, Fitch or Moody’s ratings, https://tradingeconomics.com/, or obtained by regressing OECD ratings on S&P ratings if ratings were not available.
@@
Table 4.4: Lower and upper bounds for the AVaR of the percentage losses for the real-data example discussed in Section 4.2. The underlying robust credit risk models incorporate parameter uncertainty and model uncertainty in a Gaussian, Clayton and survival Clayton copula setting as in Examples 4.1 and 4.2; see the caption of Table 4 for the respective parameters. The loss given default in this scenario is $\delta_{n}\equiv 0.1$ and the numbers are obtained by Monte Carlo simulations with a sample size of $1.000.000.$
@@
$\alpha=99$ 11.19% 11.22% 11.27% 11.36% 11.16% 11.17% 11.21% 11.32% 11.14% 11.55%
Table 4.5: Lower and upper bounds for the AVaR of the percentage losses for the real-data example discussed in Section 4.2. The underlying robust credit risk models incorporate parameter uncertainty and model uncertainty in a Gaussian, Clayton and survival Clayton copula setting as in Examples 4.1 and 4.2; see the caption of Table 4 for the respective parameters. The loss given default in this scenario is $\delta_{n}\sim beta(0.3,2.7)$-distributed and the numbers are obtained by Monte Carlo simulations with a sample size of $1.000.000.$
@@
While in the case of a homogeneous and rather large credit risk portfolio (Tables 4.1 and 4.2), the dependencies between the default events are crucial, we find that in the rather small (in terms of number of borrowers) and very heterogeneous credit portfolio of IDB, the AVaR bounds are much more comparable (Tables 4.4 and 4.5). The reason is that for such a small and concentrated (w.r.t. single borrowers) portfolio, systematic risk reflected by common dependence on the risk factor is less relevant whereas idiosyncratic risk plays a much more important role, compare Lütkebohmert et al. (2023) and Lütkebohmert and Sester (2024). Moreover, we observe that the distribution of the losses given default has a strong influence on the total portfolio risk which is due to the fact that tail events result from a combination of bad realizations of the risk factor and high values for the loss given default variable. Thus, AVaR increases with the variance of loss given default.
@@ Conclusion
In this paper, we established a convex ordering result for credit portfolio losses in stochastically increasing BMMs based on transparent and interpretable assumptions on the conditional default probabilities. For threshold models, we provided a related ordering result based on a pointwise comparison of bivariate copulas. As a consequence, we obtained several robustness results for credit portfolio losses. Without assuming monotonicity of the conditional default probabilities in the common factor, we derived an upper bound on portfolio losses that improves the standard comonotonic bound. A detailed simulation study and a real-data example illustrate the relevance and effectiveness of our approach. Our methods are especially applicable to standard industry models of credit portfolio risk, such as CreditMetrics or KMV Portfolio Manager, and allow for the incorporation of tail dependencies and model uncertainty into these models. In this way, our paper provides tractable and practically relevant methods for the assessment of credit portfolio risk under mild structural assumptions.
@@ Acknowledgements
The authors thank two anonymous referees for their valuable comments, which helped improve the manuscript. The first author gratefully acknowledges the support of the Austrian Science Fund (FWF) projects 10.55776/PAT1669224 SORT: Stochastic Orders for Functional Dependence P 36155-N ReDim: Quantifying Dependence via Dimension Reduction and the support of the WISS 2025 project ’IDA-lab Salzburg’ (20204-WISS/225/197-2019 and 20102-F1901166-KZP).
Declarations of Interest
The authors have no relevant financial or non-financial interests to declare.
Data availability
Data is publicly available via (IDB, 2022, p.83) and (S&P Global Ratings, 2022, Table 35).